OpenStudy (anonymous):

A toy company has determined that the revenue generated by a particular toy is modeled by the following equation: 13x - 0.025x^2 The variable x is measured in thousands of toys produced, and r(x) is measured in thousands of dollars. What is the maximum revenue the company can earn with this toy?

OpenStudy (elonasushchik):

Let's see the problem first: 1.) A toy company has determined that the revenue generated by a particular toy is modeled by the following equation: r(x)=13x - 0.025x² The variable x is measured in thousands of toys produced, and r(x) is measured in thousands of dollars. What is the maximum revenue the company can earn with this toy? The point at which a graph changes direction is called a maximum or a minimum. How to find the maximum (or the minimum) of a function algebraically? 1st - Take the first derivative of the function: r(x) = 13x - 0.025x² r'(x) = 13 - 2 * ( 0.025) x r'(x) = 13 - 0.05 x 2nd - Let the derivative equal zero and solve the equation this gives. There may be multiple solutions: r'(x) = 13 - 0.05x 13 - 0.05x = 0 Adding 0.05x in both terms: 13 = 0.05x Dividing both sides by 0.05: 13/0.05 = x 260 = x Since the variable x is measured in thousands of toys produced, this represents 260,000 toys. 3rd - For every solution (only one in this case), substitute x back into the original function to find the equivalent y value: r(x) = 13x - 0.025x² r(260) = 13(260) - 0.025(260)² r(260) = 1690 Since r(x) is measured in thousands of dollars, the maximum revenue the company can earn with this toy will be $ 1,690,000. ANSWER → $ 1,690,000 ------------------------

OpenStudy (elonasushchik):

Medal and fan?! Thx